OpenStudy (amonoconnor):

Another question dealing heavily with Indefinite Integrals and Anti-Derivitaives! The problem: Evaluate the following integral using U Substitution– "integral[(csc^2(2theta))(cot(2theta))(dtheta)]" ... a) solve using: "u = cot(2theta)" b) solve using: "u = csc(2theta)" Any and all help is greatly appreciated! :)

OpenStudy (jziggy):

case a) if you set u = cot(2theta) what will your du be?

OpenStudy (amonoconnor):

-csc(2theta)(dtheta)

OpenStudy (jziggy):

You need to use chainrule for that, so it'll actually be u = cot(2theta) du = -csc^2(2theta)(dtheta) So now, knowing this what will the original dtheta equal?

OpenStudy (jziggy):

du = -csc^2(2theta)(dtheta)*2, my bad :D

OpenStudy (amonoconnor):

dtheta = du/-csc^2(2theta))

OpenStudy (jziggy):

Right, so it will simplify to